OpenStudy (anonymous):

Find out the missing number 17,36,76,___,324.

OpenStudy (anonymous):

(17+1)*2=36 (36+2)*2=76 (76+3)*2=158 (158+4)*2=324

OpenStudy (anonymous):

THANX @nbouscal

OpenStudy (anonymous):

17*2+2 = 36 36*2+4 = 76 76*2+6 = 158 158*2+8 = 324